OpenStudy (anonymous):

Factor and check by multiplying: 6x + 10y -2 Not sure what are they asking?

pooja195 (pooja195):

you have to evaluate the expression

OpenStudy (anonymous):

so factor the equation first. and once you have that, multiply it out to make sure you have 6x + 10y - 2

OpenStudy (anonymous):

so what is the largest common factor of the three coefficients?

OpenStudy (anonymous):

So I came up with 2(3x +5y) is that somewhat correct?

OpenStudy (anonymous):

well, check. if you multiply that out, what do you get?

OpenStudy (anonymous):

6x + 10y

OpenStudy (anonymous):

but where I get lost is when it says -2 at the end of the equation.

OpenStudy (anonymous):

ok, what is the largest value that you can divide the smallest coefficient (2) by without getting a fraction.

OpenStudy (anonymous):

OK I dumb ok. I have not done this in 20 yrs so bear with me. So is the largest value you can divide 2

OpenStudy (anonymous):

I was not trying to insult your intelligence. I was just trying to start at the first step.

OpenStudy (anonymous):

yes, 2 is the largest value. therefore it is the greatest common factor.

OpenStudy (anonymous):

I know I apologize it just frustrates me.

OpenStudy (anonymous):

therefore, divide each of the other coefficients (6 and 10) by 2.

OpenStudy (anonymous):

So is my final answer 2(3x +5y) =6x +10y

OpenStudy (anonymous):

3/5 ?

OpenStudy (anonymous):

so the -2 is a constant, right? all you have to do is put that in the bracket as well. so you'll have 2(3x + 5y - 1) try multiplying that out.

OpenStudy (anonymous):

6x + 10y- 2

OpenStudy (anonymous):

2(3x+5y-2)

OpenStudy (anonymous):

sorry my internet was out for a couple minutes.

OpenStudy (anonymous):

no, that won't work. you'll get 6x + 10y - 4 if you have -2 inside the bracket. -2 has to be divided by 2. so inside the bracket, you should have -1.

OpenStudy (anonymous):

so if you divide each term by two you get (6x+10y-2)/2 = (3x+5y-1) so you need to multiply the entire expression (3x+5y-1) by 2. therefore it is 2(3x+5y-1)

OpenStudy (anonymous):

@bonjovifan03 have you ever looked at any of Sal Kahns videos on khanacademy.org. His math videos have helped me tremendously.

OpenStudy (anonymous):

Got it ty all.. ty tomo I will check into that... One more question I need help on and I am done for the night Simplify: 20 +4^(3 )÷(-8)

OpenStudy (anonymous):

This one i really dont get. What does the symbol mean after the four?

OpenStudy (anonymous):

follow BEDMAS ^ = exponent

OpenStudy (anonymous):

the symbol after the four is an exponent. It means you have to multiply the number preceding it that many times.

OpenStudy (anonymous):

therefore 4^3 = 4*4*4

OpenStudy (anonymous):

by order of operations you must divide 4^3 by -8 before you can add it to 20

OpenStudy (anonymous):

well I know so far its 12 divided ny -8 which doesnt work so would i find the common number they have between them?

OpenStudy (anonymous):

would the answer be 4

OpenStudy (anonymous):

why 12? 4*4 = 16*4 = 64.